OpenStudy (anonymous):

PLEASE ANSWER: f(x)= -h(x)- ((x)/(h(x))) h(-2) = -1 h ' (-2) = -1 find f ' (-2).

OpenStudy (freckles):

Hi! So do you know how to find f' first?

OpenStudy (freckles):

You need to differentiate both sides of: \[f(x)=-h(x)-\frac{x}{h(x)}\]

OpenStudy (freckles):

\[(f(x))'=f'(x)\] \[(-h(x)-\frac{x}{h(x)})'=(-h(x))'-(\frac{x}{h(x)})'=-(h(x))'-(\frac{x}{h(x)})'\] I used the constant multiple rule to bring out -1 and just look at differentiating h(x) Now I will leave the quotient rule to you: Your job is to find: \[(\frac{x}{h(x)})'\]

OpenStudy (freckles):

Let me know what you get. Just post here.

myininaya (myininaya):

Did you understand what freckles is asking?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got the answer thanks!

myininaya (myininaya):

Great! :)